The difference of the squares of two consecutive even integers is 124.
I did...
X^2-(x+2)^2=124
x^2-x^2+4x+4=124
4x=120
x=30
but she said solve it with quadratic formula... Is that even possible, if so how do I do that?
Thankyou. Sabrina (:

Answer by ankor@dixie-net.com(22740)

he difference of the squares of two consecutive even integers is 124.
:
(x+2)^2 - x^2 = 124, (x^2 is the smaller number)
FOIL (x+2)(x+2) and you have
(x^2 + 4x + 4) - x^2= 124
Remove brackets
x^2 + 4x + 4 - x^2 = 124
Combine like terms
4x + 4 = 124
You don't need the quadratic formula
4x = 124 - 4
4x = 120
x = 120/4
x = 30
:
The two integers, 30 and 32
:
Check
32^2 - 30^2
1024 - 900 = 124, confirms our solutions